OpenStudy (anonymous):

Joey's Pizza Palace sells four sizes of pizza. The mini pizza has a 6-inch diameter, and the medium pizza has a diameter twice that of the mini pizza. How many more square inches of pizza are there in the medium than in the mini pizza? Round your answer to the nearest tenth.

OpenStudy (anonymous):

Mini Pizza(diameter) = 6 inch = radius (6/2) = 3 Medium Pizza = 2wice of Minipizza(diameter) = 2*6=12 = radius = 12/2 = 6 Area of Minipizza = \[\pi2^2 = \pi4\] Area of Medium Pizza = \[\pi6^2 = \pi36\] Difference between mini and medium pizza = \[\pi36-\pi4 = \pi32 = \] (find it and round it to nearest tenth)